# Graphon Quantum Filtering systems

### v1

@@ Appendix B Link between linear and non-linear form
To apply Lemma 5 and show that the process
$W_{t}^{u}:=Y_{t}^{u}-\sqrt{\eta}\int_{0}^{t}\mathrm{t}\mathrm{r}\big{(}(L+L^{\dagger})\rho_{s,u}\big{)}\mathrm{d}s$
is a Wiener process under the new probability measure $\mathcal{P}$, it is sufficient to verify Novikov’s condition. Since $\Big{|}\mathrm{t}\mathrm{r}\big{(}(L+L^{\dagger})\rho_{s,u}\big{)}\Big{|}\leq 2\|L\|$ and $\lambda(I)=1$, then
$\mathbb{E}_{\mathcal{Q}}\big{[}\exp\{\frac{\eta}{2}\int_{0}^{T}\mathrm{t}\mathrm{r}^{2}\big{(}(L+L^{\dagger})\rho_{s,u}\big{)}\mathrm{d}s\}\big{]}$ $=\int_{I}\mathbb{E}_{\mathbb{Q}}\big{[}\exp\{\frac{\eta}{2}\int_{0}^{T}\mathrm{t}\mathrm{r}^{2}\big{(}(L+L^{\dagger})\rho_{s,u}\big{)}\mathrm{d}s\}\big{]}\lambda(\mathrm{d}u)$
$\leq\lambda(I)\exp\Big{\{}2{\eta}{}\|L\|^{2}T\Big{\}}<\infty.$
Thus, Novikov’s condition is satisfied and the Girsanov theorem applies. ∎
{acks}
[Acknowledgments]
N-H.A. and S.C. thank the Institute for Mathematical and Statistical Innovation (IMSI), supported by the National Science Foundation (Grant No. DMS-1929348), where part of this research was conducted. H.A. and G.G. acknowledge financial support by Institut Europlace de Finance.
{funding}
This work was supported by the ANR projects Q-COAST (ANR-19-CE48-0003) and IGNITION (ANR-21-CE47-0015).

### v3

@@ Appendix B Link between the linear and non-linear forms
To apply Proposition 2.8 and show that the process
$\mathbf{W}_{t}:=\mathbf{Y}_{t}-\sqrt{\eta}\int_{0}^{t}\mathrm{t}\mathrm{r}\big((L+L^{\dagger})\bar{\mathrm{R}}_{s}\big)\mathrm{d}s$
is a Wiener process under the new probability measure $\tilde{\mathcal{P}}$, it is sufficient to verify Novikov’s condition. Since $\Big|\mathrm{t}\mathrm{r}\big((L+L^{\dagger}){\rho}_{s,u}\big)\Big|\leq 2\|L\|$ and $\lambda(I)=1$, then
$\mathbb{E}_{\mathcal{Q}}\big[\exp\{\frac{\eta}{2}\int_{0}^{T}\mathrm{t}\mathrm{r}^{2}\big((L+L^{\dagger})\bar{\mathrm{R}}_{s}\big)\mathrm{d}s\}\big]$ $=\int_{I}\mathbb{E}_{\mathbb{Q}}\big[\exp\{\frac{\eta}{2}\int_{0}^{T}\mathrm{t}\mathrm{r}^{2}\big((L+L^{\dagger})\rho_{s,u}\big)\mathrm{d}s\}\big]\lambda(\mathrm{d}u)$
$\leq\exp\Big\{2{\eta}{}\|L\|^{2}T\Big\}<\infty.$
Thus, Novikov’s condition is satisfied, and hence the Girsanov theorem applies. ∎
Acknowledgments
This work was supported by the ANR projects Q-COAST (ANR-19-CE48-0003) and IGNITION (ANR-21-CE47-0015). N-H.A. and S.C. thank the Institute for Mathematical and Statistical Innovation (IMSI), supported by the National Science Foundation (Grant No. DMS-1929348), where part of this research was conducted. H. A. acknowledges support from NSF award No. 2534856. G.G. acknowledges financial support by Institut Europlace de Finance.